# Squared-field cross-correlation between kinetic Sunyaev-Zel’dovich effect and 21-cm intensity mapping

### v1

@@ II.4 RSD corrections
where $\mu=k^{\parallel}/k=\hat{n}\cdot\hat{k}$, $b_{\mathrm{HI}}$ is the bias for HI field that relates HI field with matter perturbation, and the subscript $s$ denotes for the observable in redshift space. Here we put the detailed calculation in Appendix B  [^1] For a more complete treatment of linear and non-linear RSD effect, please refer to Ref. Shaw and Lewis (2008).
Finally, we need to introduce linear bias factors to relate the perturbations in the electron density, HI distribution, and peculiar velocity field to the underlying matter density fluctuations. Since the main signal of interest comes from large-scales, the linear biasing approximation should be accurate, and we can take the electron and HI bias factors to be constants, denoted by $b_{\mathrm{e}}$ and $b_{\mathrm{HI}}$, respectively. One can also effectively define a “bias” for electron velocity field as $b_{v}\equiv aHf/(ck)$ to rewrite all the cross power spectrum in Eq. (II.3) with matter power spectrum
@@
in which the last two factors account for the RSD effect. The three terms in the first factor correspond to contributions from Terms $6$, 77 and $8$ respectively. We will show in later numerical results that, there are hierarchies among these three terms because of the differences in geometric factors in Eq. (27), which originate from the direction of velocity perturbation in Fourier space $\hat{v}(\vec{k})=\hat{k}$. Notice that here we replace the kSZ beam with a normalized filter $\tilde{f}_{\mathrm{kSZ}}$ because kSZ map is obtained from a filtered CMB map. The detailed form of $\tilde{f}_{\mathrm{kSZ}}$ is shown in the next section.
@@ III.1 Cross-correlation
We now numerically evaluate the cross-correlation function by specifying the kernel functions, beam functions and biases in Eq. (35). First, as is given in Eq. (2), the kernel function of kSZ field is just the visibility function $g(\chi)$, which can be specified for a given cosmological and astrophysical model. The kernel of HI field is given by the window function of observation
@@
where $\Delta z=z_{\mathrm{max}}-z_{\mathrm{min}}$, and $z_{\mathrm{min}}$ and $z_{\mathrm{max}}$ are the minimum and maximum redshift coverage of an observation. Based on HI observation by SKA-MID Wide Band 1 Survey Bacon et al. (2020) with approximately $20,000~\mathrm{deg}^{2}$ sky coverage and $t_{\mathrm{tot}}=10,000$ hours total integration time, we choose $z_{\mathrm{min}}=0.3$ and $z_{\mathrm{max}}=1.0$ as an estimation [^2] In fact, SKA-MID Wide Band 1 Survey planned to proceed a HI intensity mapping in a redshift range $z=0.35\sim 3$, which is wider than $z=0.3\sim 1$ we selected. Our choice allows us to perform the analysis within a region where the biases can be approximated as constant, therefore the redshift dependence of the HI bias can be neglected..
@@
Table 2: Instrumental parameters of the CMB experiments for evaluating ILC noise power. $\delta T_{\rm N}$ is the rms noise in unit of $\mu$-arcmin, which is related to the $\Delta_{T}$ quantity by $\Delta_{T}=\pi\delta T_{\rm N}/(180\times 60)$.
@@
where $\theta_{\mathrm{FWHM}}$ is the full-width at half-maximum (FWHM), with $X$ can either be HI or kSZ field. For HI detection, the FWHM is the frequency-dependent, which is
@@
where we have substituted the diameter of MeerKAT single dish $D_{\mathrm{dish}}=13.5$ m, and $\nu=1420\,{\rm MHz}/(1+z)$ is the observed frequency of $21$-cm signal at redshift $z$. One should notice, that the complete SKA-MID dish should have 133 15-meter dishes, apart from the MeerKAT dishes Bacon et al. (2020). But here we neglect the difference. For a typical redshift $z=0.5$, such a beam will lead to a $1/\mathrm{e}$ suppression factor at $\ell\simeq 110$.
For kSZ effect measurements, the instrumental parameters ACT indicates an FWHM of order $\mathcal{O}(1)$ arcmin for different frequency bands Swetz et al. (2011). Similarly, SO, as a pathfinder of CMB-S4 observation, has a similar resolution $\theta_{\mathrm{FWHM}}=1.4^{\prime}$ ($\ell_{\mathrm{max}}\simeq 8000$). Thus for a simple estimation we adopt the following uniform beam suitable for both ACT and SO
@@
Besides the beam, there is an additional factor to be applied to the observed CMB map, which is the $\ell$-space filter that suppress the primary CMB and noise components. We apply a Wiener filter  [^3] In principle, a Wiener filter should also be applied on HI data, which is neglected because we’re considering a low-noise future measurement in this work. $f(\ell)$, together with the beam $B(\ell)$, on the temperature map in harmonic space (see also, e.g. Refs. Hill et al. (2016); Ferraro et al. (2016))
@@
The total angular power spectrum $\tilde{C}_{\ell}^{\mathrm{tot}}\equiv C_{\ell}^{\mathrm{CMB}}+C_{\ell}^{\mathrm{kSZ}}+N_{\ell}$ contains contributions from primary lensed CMB temperature power spectrum $C_{\ell}^{\mathrm{CMB}}$, kSZ auto-correlation $C_{\ell}^{\mathrm{kSZ}}$, and noise angular power $N_{\ell}$. For a simple estimation, we only consider the instrumental noise with negligible foreground contamination, which is $N_{\ell}\simeq\Delta_{T}^{2}B(\ell)^{-2}$, with $\Delta_{T}^{2}$ being the rms noise per radian. A more realistic noise model accounts for residual foreground contamination, which can be estimated by evaluating the post-ILC (internal linear combination) noise that is constructed from a weighted linear combination of maps at different frequencies. We use the public package orphics444github.com/msyriac/orphics code and apply ILC method in harmonic space to obtain the noise power $N_{\ell}$ for ACT and SO respectively, whose instrument parameters are shown in Table 2 (see also ACT DR6 data release Naess et al. (2025) and SO forecast Abitbol et al. (2025)). Note that although we apply a Wiener filter to help extract the kSZ signal, our statistic may still receive contributions from CMB lensing Ferraro et al. (2016). In future work, it may be interesting to quantify the effects of CMB lensing contamination on our estimates.
@@
We finally substitute all the instrument parameters into Eq. (41) and performing the integral numerically and obtain the angular power spectrum. The different contributions received from $C^{(6)}_{\ell}$, $C^{(7)}_{\ell}$ and $C^{(8)}_{\ell}$ and the sum are shown in Fig. 2. Here the biases $b_{\mathrm{e}}$ and $b_{\mathrm{HI}}$ are set to be unity, and the contributions from each term are approximately proportional to the same constant factor $\eta=b_{\mathrm{e}}^{2}b_{\mathrm{HI}}^{2}\Omega_{\mathrm{HI}}^{2}$  [^5] Despite the non-linear effects, $C_{\ell}$ is not precisely proportional to $\eta$ because there are biases dependence in the RSD corrections.. Therefore the total cross-correlation follows a simple relationship proportional to $\eta$.
@@ III.2 SNR forecasts
Figure 3: Left–The angular power spectrum for primary (lensed) CMB (blue solid line), kSZ effect (orange solid line) and the (deconvolved) noise for ACT telescope (dashed line) and Simons Observatory (SO; dotted-dashed line). Right–The normalized filter for ACT (blue solid line) and SO (orange solid line) respectively.
@@
where $f_{\mathrm{sky}}$ is the observed sky fraction, and the tilded quantities denote for the tracers’ auto-correlations with noises. For $\tilde{C}_{\ell}^{\mathrm{kSZ}^{2}}$, we adopt Gaussian approximation Ma and Fry (2002)
@@
where $\tilde{C}_{\ell}^{\mathrm{kSZ}}=f^{2}(\ell)B^{2}(\ell)C_{\ell}^{\mathrm{tot}}$ is the kSZ-filtered CMB power spectrum. The angular power spectrum for kSZ projection field (without beam function) is given by Refs.Ma and Fry (2002); Ma and Zhao (2014)
@@
where $\mu=\hat{\ell}\cdot\hat{k}$ and $g=g(\chi)$ defined in Eq. (2). Note here without cross-correlation with HI, the integration of $\chi$ should go all the way to the comoving distance of the last scattering surface ($z_{\ast}\simeq 1100$). A good approximation is to take $\chi_{\mathrm{max}}=\chi(z=10)$, beyond which all atoms in cosmic fluid can be safely assumed to be non-ionized with $X_{\rm e}=0$. It should be noted that early-time kSZ effect (patchy reionization) would have no cross-correlation with late time HI distribution at $z\leq 1$, but can contribute to the kSZ auto-correlation. The numerical results of $_{\ell}^{\mathrm{kSZ}}, CC$$_{\ell}^{\mathrm{CMB}}$ (primary CMB) and noise level $N_{\ell}$ are show in the left panel of Fig. 3, and the normalized filters are shown in the right panel of Fig. 3. One can see from the left panel, that primary CMB is dominated the map from very large scales to $\ell\simeq 4000$, after which the kSZ becomes dominant. In addition, because the lower noise level of SO (left panel), the central kernel of the filter peaks at smaller angular scales comparing to ACT (right panel of Fig. 3).
@@
Figure 4: Power spectrum of the squared fields, in which the biases $b_{\mathrm{e}}$ and $b_{\mathrm{HI}}$ are set to be unity. Left–Power spectrum of the squared filtered kSZ field (Eq. (52)) in which both ACT (blue solid line) and SO (orange solid line) are shown. Right–Power spectrum of the squared HI IM field with (orange dashed line) and without (blue solid line) thermal noise from instruments (Eq. (LABEL:eq:_Cl_HI_PS_field)).
@@
Figure 5: Signal-to-Noise Ratio (SNR) per-$\ell$ mode ($\mathrm{SNR}_{\ell}$, left) and cumulative SNR (SNRc, right) as functions of $\ell_{\mathrm{max}}$ for different CMB observations cross-correlating with SKA-MID, where in the right panel the horizontal dashed gray line denotes the benchmark ${\rm SNR}_{\rm c}=3$.
@@
where $P_{\mathrm{noise}}(k)$ is the 1-D noise power spectrum for HI observation. Although both $\tilde{C}_{\ell}^{\mathrm{kSZ}^{2}}$ and $\tilde{C}_{\ell}^{\mathrm{HI}^{2}}$ contain convolutions of the power spectrum, they behave differently because kSZ field is a projection field with no parallel mode information, but HI field is not. The noise mainly consists of the thermal noise from instruments, i.e. $P_{\mathrm{noise}}=P_{\mathrm{T}}$, whose power spectrum can be evaluated as Bull et al. (2015); Jiang et al. (2023)
@@
where the overall system temperature would be $T_{\mathrm{sys}}\simeq 22~\mathrm{K}$ (at $1200~\mathrm{MHz}$Bacon et al. (2020), $A_{\rm e}$ is the effective collecting area of a dish, $A_{\rm s}$ is the survey area, and $V_{\mathrm{voxel}}=(\chi\theta_{\mathrm{b}})^{2}\delta\chi$ is the comoving volume of each voxel. Here $\theta_{\mathrm{b}}$ is the FWHM of the antenna and comoving distance resolution $\delta\chi$ at given redshift is evaluated via Eq. (69) with frequency resolution $\delta\nu=0.2~\mathrm{MHz}$. The total observation time $t_{\mathrm{tot}}=10^{4}\,{\rm hrs}$ for each dish leads to a relatively small enough noise power $P_{\mathrm{T}}\sim\mathcal{O}(1)~\mu\mathrm{K}^{2}$ that can be neglected comparing with the HI signal at large scale, suggesting a good approximation $\tilde{P}_{\mathrm{HI}}(k)\simeq P_{\mathrm{HI}}(k)$ for $\ell\leq 200$. The power spectra of kSZ square field and 21-cm-squared field are show in the left and right panels of Fig. 4 respectively.
@@
One of the most important factors that affect the SNR is the thermal noise of HI observation. As is shown in the left panel of Fig. 5, the SNR per-$\ell$ begins to drop rapidly at the scale where thermal noise begins to dominate the auto-correlation of HI square field. It should be noted that it is not the total observation time but the time per area that affects the noise power, as a proportion relationship $P_{\mathrm{T}}\propto A_{\mathrm{s}}/t_{\mathrm{tot}}$ is suggested from the definition Eq. (56). For simplicity let us assume the $10,000$ hours integration time mentioned is uniformly distributed into $20,000~\mathrm{deg}^{2}$. As HI signal is going to cross-correlate with the kSZ observation with an overlap sky-coverage $8,000~\mathrm{deg}^{2}$, the effective integration time is actually $4,000$ hours.
At last, here we provide a rough estimation on the results with a shorter integration time, for example $1,000$ hours for $8,000~\mathrm{deg}^{2}$ sky-coverage, in which case the thermal noise would begin to dominate HI square field auto-signal at a larger scale $\ell\simeq 100$, leading to a suppression of SNR per mode starting at a smaller $\ell$. Then the approximate cumulative SNR would reach its maximum at an earlier stage. Let us truncate the cumulation at $\ell\simeq 100$, then the results finally arrive at $\mathrm{SNR}_{\mathrm{max}}\simeq 3$ for SO and $\mathrm{SNR}_{\mathrm{max}}\simeq 1.5$ for ACT.

### v2

@@ II.4 Redshift-space distortion (RSD) corrections
where $\mu=k^{\parallel}/k=\hat{n}\cdot\hat{k}$, $b_{\mathrm{HI}}$ is the bias for the HI field that relates the HI field with matter perturbation, and the subscript $s$ denotes for the observable in redshift space. Here we put the detailed calculation in Appendix B  [^1] For a more complete treatment of linear and nonlinear RSD effects, please refer to Ref. Shaw and Lewis (2008).
Finally, we need to introduce linear bias factors to relate the perturbations in the electron density, HI distribution, and peculiar velocity field to the underlying matter density fluctuations. Since the main signal of interest comes from large scales, the linear biasing approximation should be accurate, and we can take the electron and HI bias factors to be constants, denoted by $b_{\mathrm{e}}$ and $b_{\mathrm{HI}}$, respectively. One can also effectively define a “bias” for electron velocity field as $b_{v}\equiv aHf/(ck)$ to rewrite all the cross power spectrum in Eq. (II.3) with matter power spectrum
@@
in which the last two factors account for the RSD effect. The three terms in the first factor correspond to contributions from Terms $6$, 77 and $8$ respectively. We will show in later numerical results that, there are hierarchies among these three terms because of the differences in geometric factors in Eq. (27), which originate from the direction of velocity perturbation in Fourier space $\hat{v}(\vec{k})=\hat{k}$. Notice that here we replace the kSZ beam with a normalized filter $\tilde{f}_{\mathrm{kSZ}}$ because the kSZ map is obtained from a filtered CMB map. The detailed form of $\tilde{f}_{\mathrm{kSZ}}$ is shown in the next section.
@@ III.1 Cross-correlation
We now numerically evaluate the cross-correlation function by specifying the kernel functions, beam functions and biases in Eq. (35). First, as is given in Eq. (2), the kernel function of the kSZ field is just the visibility function $g(\chi)$, which can be specified for a given cosmological and astrophysical model. The kernel of the HI field is given by the window function of observation
@@
where $\Delta z=z_{\mathrm{max}}-z_{\mathrm{min}}$, and $z_{\mathrm{min}}$ and $z_{\mathrm{max}}$ are the minimum and maximum redshift coverage of an observation. Based on HI observation by SKA-MID Wide Band 1 Survey Bacon et al. (2020) with approximately $20,000~\mathrm{deg}^{2}$ sky coverage and $t_{\mathrm{tot}}=10,000$ hours total integration time, we choose $z_{\mathrm{min}}=0.3$ and $z_{\mathrm{max}}=1.0$ as an estimation. [^2] In fact, SKA-MID Wide Band 1 Survey planned to proceed a HI intensity mapping in a redshift range $z=0.35\sim 3$, which is wider than $z=0.3\sim 1$ we selected. Our choice allows us to perform the analysis within a region where the biases can be approximated as constant, therefore the redshift dependence of the HI bias can be neglected.
@@
Table 2: Instrumental parameters of the CMB experiments for evaluating internal linear combination (ILC) noise power. $\delta T_{\rm N}$ is the rms noise in unit of $\mu$-arcmin, which is related to the $\Delta_{T}$ quantity by $\Delta_{T}=\pi\delta T_{\rm N}/(180\times 60)$.
@@
where $\theta_{\mathrm{FWHM}}$ is the full-width at half-maximum (FWHM), with $X$ can be either HI or kSZ field. For HI detection, the FWHM is frequency dependent, which is
@@
where we have substituted the diameter of MeerKAT single dish $D_{\mathrm{dish}}=13.5$ m, and $\nu=1420\,{\rm MHz}/(1+z)$ is the observed frequency of the $21$-cm signal at redshift $z$. One should notice, that the complete SKA-MID dish should have 133 15-meter dishes, apart from the MeerKAT dishes Bacon et al. (2020). But here we neglect the difference. For a typical redshift $z=0.5$, such a beam will lead to a $1/\mathrm{e}$ suppression factor at $\ell\simeq 110$.
For the kSZ effect measurements, the instrumental parameters ACT indicate an FWHM of order $\mathcal{O}(1)$ arcmin for different frequency bands Swetz et al. (2011). Similarly, SO, as a pathfinder of CMB-S4 observation, has a similar resolution $\theta_{\mathrm{FWHM}}=1.4^{\prime}$ ($\ell_{\mathrm{max}}\simeq 8000$). Thus for a simple estimation we adopt the following uniform beam suitable for both ACT and SO
@@
Besides the beam, there is an additional factor to be applied to the observed CMB map, which is the $\ell$-space filter that suppress the primary CMB and noise components. We apply a Wiener filter  [^3] In principle, a Wiener filter should also be applied on HI data, which is neglected because we’re considering a low-noise future measurement in this work. $f(\ell)$, together with the beam $B(\ell)$, on the temperature map in harmonic space (see also, e.g., Refs. Hill et al. (2016); Ferraro et al. (2016))
@@
The total angular power spectrum $\tilde{C}_{\ell}^{\mathrm{tot}}\equiv C_{\ell}^{\mathrm{CMB}}+C_{\ell}^{\mathrm{kSZ}}+N_{\ell}$ contains contributions from primary lensed CMB temperature power spectrum $C_{\ell}^{\mathrm{CMB}}$, kSZ autocorrelation $C_{\ell}^{\mathrm{kSZ}}$, and noise angular power $N_{\ell}$. For a simple estimation, we only consider the instrumental noise with negligible foreground contamination, which is $N_{\ell}\simeq\Delta_{T}^{2}B(\ell)^{-2}$, with $\Delta_{T}^{2}$ being the rms noise per radian. A more realistic noise model accounts for residual foreground contamination, which can be estimated by evaluating the post-ILC (internal linear combination) noise that is constructed from a weighted linear combination of maps at different frequencies. We use the public package orphics444github.com/msyriac/orphics code and apply the ILC method in harmonic space to obtain the noise power $N_{\ell}$ for ACT and SO respectively, whose instrument parameters are shown in Table 2 (see also ACT DR6 data release Naess et al. (2025) and SO forecast Abitbol et al. (2025)). Note that although we apply a Wiener filter to help extract the kSZ signal, our statistic may still receive contributions from CMB lensing Ferraro et al. (2016). In future work, it may be interesting to quantify the effects of CMB lensing contamination on our estimates.
@@
We finally substitute all the instrument parameters into Eq. (41), perform the integral numerically and obtain the angular power spectrum. The different contributions received from $C^{(6)}_{\ell}$, $C^{(7)}_{\ell}$ and $C^{(8)}_{\ell}$ and the sum are shown in Fig. 2. Here the biases $b_{\mathrm{e}}$ and $b_{\mathrm{HI}}$ are set to be unity, and the contributions from each term are approximately proportional to the same constant factor $\eta=b_{\mathrm{e}}^{2}b_{\mathrm{HI}}^{2}\Omega_{\mathrm{HI}}^{2}$. [^5] Despite the nonlinear effects, $C_{\ell}$ is not precisely proportional to $\eta$ because there are biases dependence in the RSD corrections. Therefore the total cross-correlation follows a simple relationship proportional to $\eta$.
@@ III.2 SNR forecasts
Figure 3: Left: The angular power spectrum for primary (lensed) CMB (blue solid line), kSZ effect (orange solid line) and the (deconvolved) noise for the ACT telescope (dashed line) and Simons Observatory (SO; dot-dashed line). Right: The normalized filter for ACT (blue solid line) and SO (orange solid line) respectively.
@@
where $f_{\mathrm{sky}}$ is the observed sky fraction and the tilded quantities denote the tracers’ autocorrelations with noises. For $\tilde{C}_{\ell}^{\mathrm{kSZ}^{2}}$, we adopt Gaussian approximation Ma and Fry (2002)
@@
where $\tilde{C}_{\ell}^{\mathrm{kSZ}}=f^{2}(\ell)B^{2}(\ell)C_{\ell}^{\mathrm{tot}}$ is the kSZ-filtered CMB power spectrum. The angular power spectrum for the kSZ projection field (without beam function) is given by Refs.Ma and Fry (2002); Ma and Zhao (2014)
@@
where $\mu=\hat{\ell}\cdot\hat{k}$ and $g=g(\chi)$ are defined in Eq. (2). Note here without cross-correlation with HI, the integration of $\chi$ should go all the way to the comoving distance of the last scattering surface ($z_{\ast}\simeq 1100$). A good approximation is to take $\chi_{\mathrm{max}}=\chi(z=10)$, beyond which all atoms in cosmic fluid can be safely assumed to be nonionized with $X_{\rm e}=0$. It should be noted that the early-time kSZ effect (patchy reionization) would have no cross-correlation with late time HI distribution at $z\leq 1$, but can contribute to the kSZ autocorrelation. The numerical results of $_{\ell}^{\mathrm{kSZ}}, CC$$_{\ell}^{\mathrm{CMB}}$ (primary CMB) and noise level $N_{\ell}$ are show in the left panel of Fig. 3, and the normalized filters are shown in the right panel of Fig. 3. One can see from the left panel that primary CMB dominates the map from very large scales to $\ell\simeq 4000$, after which the kSZ becomes dominant. In addition, because of the lower noise level of SO (left panel), the central kernel of the filter peaks at smaller angular scales comparing to ACT (right panel of Fig. 3).
@@
Figure 4: Power spectrum of the squared fields, in which the biases $b_{\mathrm{e}}$ and $b_{\mathrm{HI}}$ are set to be unity. Left: Power spectrum of the squared filtered kSZ field (Eq. (52)) in which both ACT (blue solid line) and SO (orange solid line) are shown. Right: Power spectrum of the squared HI IM field with (orange dashed line) and without (blue solid line) thermal noise from instruments (Eq. (LABEL:eq:_Cl_HI_PS_field)).
@@
Figure 5: Signal-to-noise ratio per-$\ell$ mode ($\mathrm{SNR}_{\ell}$, left) and cumulative SNR (SNRc, right) as functions of $\ell_{\mathrm{max}}$ for different CMB observations cross-correlating with SKA-MID, where in the right panel the horizontal dashed gray line denotes the benchmark ${\rm SNR}_{\rm c}=3$.
@@
where $P_{\mathrm{noise}}(k)$ is the 1D noise power spectrum for HI observation. Although both $\tilde{C}_{\ell}^{\mathrm{kSZ}^{2}}$ and $\tilde{C}_{\ell}^{\mathrm{HI}^{2}}$ contain convolutions of the power spectrum, they behave differently because the kSZ field is a projection field with no parallel mode information, but HI field is not. The noise mainly consists of the thermal noise from instruments, i.e. $P_{\mathrm{noise}}=P_{\mathrm{T}}$, whose power spectrum can be evaluated as Bull et al. (2015); Jiang et al. (2023)
@@
where the overall system temperature would be $T_{\mathrm{sys}}\simeq 22~\mathrm{K}$ (at $1200~\mathrm{MHz}$Bacon et al. (2020), $A_{\rm e}$ is the effective collecting area of a dish, $A_{\rm s}$ is the survey area, and $V_{\mathrm{voxel}}=(\chi\theta_{\mathrm{b}})^{2}\delta\chi$ is the comoving volume of each voxel. Here $\theta_{\mathrm{b}}$ is the FWHM of the antenna and comoving distance resolution $\delta\chi$ at a given redshift is evaluated via Eq. (69) with frequency resolution $\delta\nu=0.2~\mathrm{MHz}$. The total observation time $t_{\mathrm{tot}}=10^{4}\,{\rm hrs}$ for each dish leads to a relatively small enough noise power $P_{\mathrm{T}}\sim\mathcal{O}(1)~\mu\mathrm{K}^{2}$ that can be neglected comparing with the HI signal at large scale, suggesting a good approximation $\tilde{P}_{\mathrm{HI}}(k)\simeq P_{\mathrm{HI}}(k)$ for $\ell\leq 200$. The power spectra of the kSZ square field and 21-cm-squared field are show in the left and right panels of Fig. 4 respectively.
@@
One of the most important factors that affect the SNR is the thermal noise of the HI observation. As is shown in the left panel of Fig. 5, the SNR per-$\ell$ begins to drop rapidly at the scale where thermal noise begins to dominate the autocorrelation of the HI square field. It should be noted that it is not the total observation time but the time per area that affects the noise power, as a proportion relationship $P_{\mathrm{T}}\propto A_{\mathrm{s}}/t_{\mathrm{tot}}$ is suggested from the definition Eq. (56). For simplicity let us assume the $10,000$ hours integration time mentioned is uniformly distributed into $20,000~\mathrm{deg}^{2}$. As the HI signal is going to cross-correlate with the kSZ observation with an overlap sky-coverage $8,000~\mathrm{deg}^{2}$, the effective integration time is actually $4,000$ hours.
At last, here we provide a rough estimation on the results with a shorter integration time, for example $1,000$ hours for $8,000~\mathrm{deg}^{2}$ sky-coverage, in which case the thermal noise would begin to dominate the HI square field autosignal at a larger scale $\ell\simeq 100$, leading to a suppression of SNR per mode starting at a smaller $\ell$. Then the approximate cumulative SNR would reach its maximum at an earlier stage. Let us truncate the cumulation at $\ell\simeq 100$, and then the results finally arrive at $\mathrm{SNR}_{\mathrm{max}}\simeq 3$ for SO and $\mathrm{SNR}_{\mathrm{max}}\simeq 1.5$ for ACT.